OpenStudy (anonymous):

PLEASE HELP A ball of mass 0.250 kg on the end of 0.850 m long cord (ignorable mass) iS swung in a vertical circle. Determine: (a) the minimum speed the ball must have at the top of its arc so that it continues moving in a circle (b) the tension in the cord at the bottom of the arc if the ball is moving at twice that speed

OpenStudy (ulahlynn):

hi my sister! :D

OpenStudy (jfraser):

i assume you mean for part a) what velocity it must have to keep moving in a circle. What you need to see is that the KE of the ball MUST EQUAL the PE of its height in order to keep moving in the circular path. Set KE = PE\[\frac{1}{2}m*v^2 = m*g*h\]and solve for v

OpenStudy (anonymous):

oh okay, I get it. how about the tension? do we need to get the normal force and stuff?

OpenStudy (anonymous):

:)

OpenStudy (jfraser):

there is no normal force because the ball isn't in contact with a surface like a table. You need to find the centripetal force, which is equal to the tension in the rope. Using the velocity you found from part a), plug it into a formula so you can find F. Something like:\[F_c = m*a_c\]but since \(a_c\) is the centripetal acceleration, we can substitute and get \[F_c = \frac{m*v^2}{r}\]

OpenStudy (anonymous):

oh sorry bout that. thankyou very much

OpenStudy (anonymous):

can I ask another one? :)

OpenStudy (jfraser):

no need to apologize, just have to pay attention to what all the formulas in physics mean

OpenStudy (jfraser):

sure. close this and ask another